OpenStudy (anonymous):

I tried to find this out but failed six over seven k = 6 what does k= in this equation example > k=5 (not true but an exsample)

OpenStudy (anonymous):

@Compassionate

OpenStudy (compassionate):

What? Lol.

OpenStudy (anonymous):

6 over 7 k, whats k equal in this equation XD

OpenStudy (compassionate):

I'm not an administrator. I cannot login to that course.

OpenStudy (compassionate):

\[\frac{ 6 }{ 7k} = 6\] Just solve the equation for k. Since 6 is being divided, multiply both sides by 6.

OpenStudy (anonymous):

calculator was playing games on me T.T

OpenStudy (anonymous):

@Compassionate gave you a medal and everytime I tried I dunno why but it came up as 0.14285714285 on calculator I knew it wouldave been 6 or 7 but I just got conffizled thankies ^^

OpenStudy (compassionate):

No, you're solving for k.

OpenStudy (anonymous):

I know >w< I was saying what k was possibly going to equal how it showed me something different on calculator

OpenStudy (ranga):

Multiply both sides by k. Divide both sides by 6. and you get k = 1/7 (you can leave it as a fraction. But in decimal, 1/7 = 0.1428571428)

OpenStudy (compassionate):

\[k = \frac{ 1 }{ 7 }\]